Writing Equations
The slope is 1; the y intercept is 2

Answer by unlockmath(1688)

Hello,
Here it is:
y=1x+2 or written as y=x+2
We know the slope-intercept form is:
y=mx+b
M is the slope
b is where the line crosses the y axis.
Make sense?
RJ
www.math-unlock.com